OpenStudy (anonymous):

Did I do it right? The range of the function f(x) = x + 5 is {7, 9}. What is the function’s domain? {2, 4} {-2, -4} {12, 14} {-12, -14} {0, 5} A? Just checking

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

yep, 2+5 and 4+5 get us 7 and 9

OpenStudy (anonymous):

Okay just wanted to make sure that 7&9 go where f(x)

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

x+5 = 7,9 -5 -5 x = 2,4